A shipment of 20 similar laptop computers to a retail outlet contains 3 that are defective. If a school makes a random purchase of 2 of these computers, find the probability distribution for the number of defectives. Show the table in your solution. Round off answers to 4 decimal places.


1
Expert's answer
2022-01-19T03:22:38-0500

Let "X=" the number of defective computers purchased: "X=0,1,2."


"P(X=0)=\dfrac{\dbinom{3}{0}\dbinom{20-3}{2-0}}{\dbinom{20}{2}}"

"=\dfrac{1(136)}{190}=\dfrac{68}{95}"


"P(X=1)=\dfrac{\dbinom{3}{1}\dbinom{20-3}{2-1}}{\dbinom{20}{2}}"

"=\dfrac{3(17)}{190}=\dfrac{51}{190}"



"P(X=2)=\dfrac{\dbinom{3}{2}\dbinom{20-3}{2-2}}{\dbinom{20}{2}}"

"=\dfrac{3(1)}{190}=\dfrac{3}{190}"

The probability distribution for the number of defectives


"\def\arraystretch{1.5}
 \begin{array}{c:c}
 x & 0 & 1 & 2 \\ \hline
 p(x) & 0.7158 & 0.2684 & 0.0158 \\
 
\end{array}"